OpenStudy (anonymous):

PLEASE HELP! LAST DAY TO TURN IN WORK FOR GRADUATION! If f(x) = 4x + 1 and g(x) = x2 – 2x + 1, what is f(g(x)) ?

OpenStudy (across):

You don't deserve to graduate! :D I'm just kidding. Do you really not know how to perform function composition? Let\[f(x)=4x+1\]and\[g(x)=x^2-2x-1\]Then\[f(g(x))=4(x^2-2x-1)+1\]Can you simplify that? You better.

OpenStudy (across):

I made a typo on purpose. ;P You better find it and fix it, too.

OpenStudy (anonymous):

i honestly wish i did know how to do all of this work, but since my teacher did know half the crap she was doing, she ended up making the whole class fail....thanks for helping though

OpenStudy (across):

I mean, it's not like these are nuclear physics concepts either. A high school student could literally pass every single course with an A+ if this just sat down and read the book for 10 minutes every day.

OpenStudy (anonymous):

not really, im ADD.

OpenStudy (anonymous):

perhaps leaving this until the last day wasn't the best decision...

OpenStudy (anonymous):

what's ADD?

OpenStudy (across):

ADD + procrastination I suggest you work on that asap.

OpenStudy (amistre64):

f(x) is a generic way of naming a function; the rule for the function is defined by an equation that has an "x" in it. when the inside of it is replaced by something else, it means that x take on the value of that something else. f(g(x)) means that the x part of f(x) is replaced by the rule or definition of g(x)

OpenStudy (anonymous):

i have 3 assignments left, the rest of the work i struggled to do myself. these are the 3 i have trouble with. and i think i did pretty good considering i have to work in order to make sure me and my mom can keep a roof over our heads.

OpenStudy (amistre64):

f(x) = 4x + 1 f(g(x)) = 4(g(x)) + 1 ; and g(x) = x^2 – 2x + 1

OpenStudy (amistre64):

your best bet is to ask us what about this you might not understand so that we can give you a simpler picture of it